Can you please help show me how to solve this quadratic equation by completing the square.
x^2 + 4x - 12 = 0

Answer by stanbon(75887)

solve this quadratic equation by completing the square.
x^2 + 4x - 12 = 0
----
x^2 + 4x + ? = 12 + ?
x^2 + 4x + (4/2)^2 = 12 + (4/2)^2
x^2 + 4x + 4 = 12 + 4
(x+2)^2 = 16
x+2 = 4 or x+2 = -4
x = 2 or x = -6
=====================
Cheers,
Stan H.